OpenStudy (anonymous):

LEAST TO GREATEST! 2/16, 2/4, 9/10, -0.078! Thanks!

OpenStudy (anonymous):

-.078,2/16,2/4,9/10

OpenStudy (anonymous):

2/16 = 1/8 or .125 2/4 = 1/2 or .5 9/10 or .9 so from least to greatest: -0.078, 2/16, 2/4, 9/10

OpenStudy (anonymous):

Thank you guys